OpenStudy (jenniferjuice):

What are the solutions of x2 + 5x − 24 = 0? Show all work for full credit.

OpenStudy (anonymous):

use the quadratic equation on your 2nd order polynomial that is in the form of\[ax^2 + bx + c \] finding both x values, using the plus or minus sign: \[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\]

hero (hero):

Or you could just factor \[x^2 + 5x - 24 = 0\]\[x^2 +8x - 3x -3(8) = 0\]\[x(x + 8) - 3(x + 8) = 0\]\[(x + 8)(x - 3) = 0\]Use zero product property to finish solving.

OpenStudy (anonymous):

use Euler271's suggestion as a last resort when you can't factor it out. The problem I have with it is that I can be careless sometimes and leave out a "-" or "+" and it will give a wrong solution. boths solutions are good...I just prefer to factor it out first then use the quadratic equation if I can't solve it.